please help me with these problems.
Write the equation of the line passing through each of the given pairs, write your result in slope-intercept form, when possible.
(-1,3)and (4,-2)
(2,-3)and (2,4)
Thank you for your help

Answer by checkley75(3666)

(-1,-3) & (4,-2) the slope is (-2+3)/(4+1)=1/5. now we find the y intercept (b)
Y=mX+b BY SUBSTITUTING ONE SET OF POINTS FOR X & Y THUS:
3=1/5*-1+b
3=-1/5+b
b=3+1/5
b=16/5
SO THIS LINE EQUATION IS: Y=X/5+16/5
-----------------------------------------------------------------
(2,-3) & (2,4) THE SLOPE IS (4+3)/(2-2)=7/0 WHICH IS A VERTICAL LINE THROUGH THE POINT X=2
THUS THE LINE EQUATION HERE IS X=2. NO Y INTERCEPT.